OpenStudy (anonymous):

Find the vertex of the graph of the function. f(x) = 4x^2 + 24x + 32

OpenStudy (anonymous):

how do i put it in vertex form

OpenStudy (anonymous):

you need to complete the square. first take out 4 as a common factor: 4x^2 + 24x + 32 = 4(x^2 + 6x + 8) 6/2 = 3 so: = 4( (x+3)^2 - 3^2 + 8 ) = 4( (x+3)^2 - 9 + 8) = 4( (x+3)^2 - 1) = 4(x+3)^2 - 4 vertex will be the minimum point, so: now, 4(x+3)^2 - 4 is at a minimum when x = -3 because it makes the stuff being squared equal to 0 so we know the x coord of vertex is -3 the y coord is the constant on the end, -4

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

sorry ill explain a bit more :) anything that confuses you dont be worried about saying. im not very good at explaining so it'll probably be my fault

OpenStudy (anonymous):

4x^2 + 24x + 32 the bit about the 4 being a common factor isnt that important, but it makes things a bit easier. i realized it was a common factor because 4x^2 is 4*(x^2) , 24x is 4*(6x) and 32 is 4*(8) can you see that they are all multiples of 4? don't worry if you don't or if you take a lot of time to, it all comes with experience.

OpenStudy (anonymous):

oh i seee.... okay and then next you put 4(x+3)^2-3^2+8 can you explain what you did in the next 3 steps

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

4x^2 + 24x + 32 = 4(x^2 + 6x + 8) lets just forget about the 4 for a minute, and we'll remember him later we want to "complete the square" for: x^2 + 6x + 8 the basic idea is this look at this: (x+3)^2 = (x+3)(x+3) = x^2 + 3x + 3x + 9 = x^2 + 6x + 9 this is almost what we have isn't it?

OpenStudy (anonymous):

Are you real teachers?

OpenStudy (anonymous):

if (x+3)^2 = x^2 + 6x + 9 then (x + 3)^2 - 1 = (x^2 + 6x + 9) -1 = x^2 + 6x + 8 so (x + 3)^2 - 1 = x^2 + 6x + 8 the reason (x+3) works is because 3 is half of 6

OpenStudy (anonymous):

@headge im not a teacher, im a student

OpenStudy (anonymous):

WOOOOOW

OpenStudy (anonymous):

Eigenschmeigen?

OpenStudy (anonymous):

@eigenschmeigen ?

OpenStudy (anonymous):

?